OpenStudy (deez):

Find the General solution for the following ODE: (x^3 + y\x).dx + (y^2 + ln (x) ).dy = 0

OpenStudy (anonymous):

x^3dx +ydx/x + y^2dy + lnxdy = 0 x^3dx + y^2dy + (ydx/x + lnxdy) = 0 d(ylnx) = -[x^4 / 4 + y^3 / 3] ylnx = -[x^4 / 4 + y^3 / 3]

OpenStudy (lalaly):

haraaam lssa bta5od el araf hada....hay azn5 madeh ba5odha...

OpenStudy (anonymous):

? lol

OpenStudy (lalaly):

hahahahahaha